OpenStudy (anonymous):

what is the relative maximum of f(x)=−3x^2+5x+2 (the relative minimum does not exist) please help!!!!

OpenStudy (anonymous):

differentiate it and equate to zero

OpenStudy (anonymous):

you will get the value of x where it is maximum and that is 5/6

OpenStudy (anonymous):

Substitute the value of X in f(x)

OpenStudy (anonymous):

plug in -2 for every x within the equation f(x) so f(-2)=3 (-2^2) + 5 (-2) -2 =3 (4) +(-10) -2 =12-10-2 = 0

OpenStudy (anonymous):

Wait so what does that mean? It says the answer is supposed to be (x,y)

OpenStudy (anonymous):

is it (5/6,49/12)?

OpenStudy (anonymous):

No :/